OpenStudy (anonymous):

what is the 5th partial sum of infinity *sigma notation* -3+5n n=1

OpenStudy (anonymous):

Answer choices: 55 60 65 70 Please expain it too, cause I would like to know how to do another one

OpenStudy (zzr0ck3r):

1st term, plug in 1 to -3+5n and get 2 2nd term put in two and get 7 .... 5ith term put in 5 add them all up.

OpenStudy (anonymous):

how high would I go though? Cause it says infinity

OpenStudy (zzr0ck3r):

2+7+12+17+22= 60

OpenStudy (zzr0ck3r):

5 terms

OpenStudy (zzr0ck3r):

the series is defined for infinity but when they ask for the nth partial they only want the first n terms

OpenStudy (anonymous):

ohh so if they want the 5th partial sum, you add up the first 5 terms?

OpenStudy (amistre64):

i simpler way to write that would be sum (1 to inf) of: -3+5n

OpenStudy (zzr0ck3r):

I dont think it is phrazed right, its the 5th partial sum of sumation blah blah from to infinity

OpenStudy (zzr0ck3r):

correct mother

OpenStudy (zzr0ck3r):

its really sum (1 to 5) of: -3+5n

OpenStudy (amistre64):

does the rule define a sequence that then gets partially sumed?

OpenStudy (zzr0ck3r):

\[\sum_{1}^{5} -3+5n\]

OpenStudy (anonymous):

so wait say in this problem: \[\sum_{8}^{n=5}\]

OpenStudy (zzr0ck3r):

yeah I think so but Im just saying the computation is like 1->5

OpenStudy (anonymous):

a (with a sub-note) of n next to sigma notation

OpenStudy (amistre64):

\[\sum_{1}^{5} -3+5n\] \[\sum_{1}^{5} -3+\sum_{1}^{5} 5n\] \[ -3\sum_{1}^{5}+5\sum_{1}^{5}n\]

OpenStudy (zzr0ck3r):

8->5 does not make any sense

OpenStudy (anonymous):

hold on. my computer is going slow so when it was loading it switched the numbers

OpenStudy (zzr0ck3r):

but even if you go "backwards" you will get the same thing so...

OpenStudy (anonymous):

8 sigma notation a (sub-note) n n=5

OpenStudy (anonymous):

thats it^ my computer wasnt loading right.

OpenStudy (anonymous):

how do you put that in a calculator?

OpenStudy (amistre64):

internet explorer is a bad browser for this site; if you havent already, then you might wanna switch to the goolge chrome for all around internetting performance improvements

OpenStudy (zzr0ck3r):

\[\sum_{5}^{8}-3+2n\] = (-3+2(5))+(-3+2(6))+(-3+2(7))+(-3+2(8))

OpenStudy (anonymous):

no theres a(n) next to the sigma notation. But the n is like below it

OpenStudy (zzr0ck3r):

same thing

OpenStudy (zzr0ck3r):

yeah same thing

OpenStudy (zzr0ck3r):

the way you wrote is the better notation...

OpenStudy (anonymous):

but how you put that in a calculator?

OpenStudy (anonymous):

causse the n is below it

OpenStudy (zzr0ck3r):

depends on the calculator

OpenStudy (zzr0ck3r):

for ti89 i think its sigma(-3+2n, n, 5, 8)

OpenStudy (anonymous):

is there anyway to do it online?

OpenStudy (zzr0ck3r):

the n is jsut telling you the variable name so like \[\sum_{n=1}^{8}3x\] would not make sense to talk about.

OpenStudy (anonymous):

but wait theres 2 variables a and n. Im sorry Im completely confused on this one

OpenStudy (zzr0ck3r):

A is the function name A(n) = A_n = -3+2n for natural numbers n this is the notation we use to show we are talking about n as integers and not real numbers

jimthompson5910 (jim_thompson5910):

there is a way to do it online if you type sum(-3+5n,n=1..5) into wolfram alpha, you'll get \[\Large \sum_{n=1}^{5} -3+5n = 60\] wolfram alpha: www.wolframalpha.com

jimthompson5910 (jim_thompson5910):

of course, this should be a way to check your work (and not do it completely for you)

OpenStudy (zzr0ck3r):

also get to know the such that key | on your calculator, it will make life easy

OpenStudy (anonymous):

thankyou! can you help me with the one below

OpenStudy (anonymous):

Im just really confused on that one still

jimthompson5910 (jim_thompson5910):

\[\Large \sum_{n=5}^{8} a_{n}\] literally means "add the terms a5, a6, a7, a8" So \[\Large \sum_{n=5}^{8} a_{n} = a_{5}+a_{6}+a_{7}+a_{8}\] The only problem is that I don't know what \(\Large a_{n}\) is (is it \(\Large a_{n} = -3+5n\) ?)

OpenStudy (anonymous):

ohh cause my answer choices are 54 102 144 162

jimthompson5910 (jim_thompson5910):

what is the sequence \(\Large a_{n}\) in this case? You didn't specify

OpenStudy (anonymous):

-3,0,3

jimthompson5910 (jim_thompson5910):

so it looks like you're adding 3 each time

jimthompson5910 (jim_thompson5910):

those are the first three terms, what are the next 5 terms?

OpenStudy (anonymous):

6, 9, 12, 15, 18,

jimthompson5910 (jim_thompson5910):

Now list the 5th, 6th, 7th, and 8th terms

OpenStudy (anonymous):

18, 21, 24, 27

jimthompson5910 (jim_thompson5910):

Sorry, I should have stated that the first 8 terms are... -3, 0, 3, 6, 9, 12, 15, 18 This is of course assuming that -3, 0, 3 are the first three terms (that are given)

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

so the 5th, 6th, 7th, and 8th terms of the sequence -3, 0, 3, 6, 9, 12, 15, 18 are...???

OpenStudy (anonymous):

21,24,27,30

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

they're in that list somewhere

OpenStudy (anonymous):

sorry that wasnt the 5th, 6th 7th. I didnt read that part

jimthompson5910 (jim_thompson5910):

thats ok

OpenStudy (anonymous):

9, 12, 15, 18

jimthompson5910 (jim_thompson5910):

now add those terms up

OpenStudy (anonymous):

54

OpenStudy (anonymous):

thats the answer?!

jimthompson5910 (jim_thompson5910):

good, so... \[\Large \sum_{n=5}^{8} a_{n} = a_{5}+a_{6}+a_{7}+a_{8}\] \[\Large \sum_{n=5}^{8} a_{n} = 9+12+15+18\] \[\Large \sum_{n=5}^{8} a_{n} = 54\]

jimthompson5910 (jim_thompson5910):

it is

OpenStudy (anonymous):

thankyou SO much!

jimthompson5910 (jim_thompson5910):

you're welcome